OpenStudy (anonymous):

integrate csc^5 x dx

OpenStudy (aravindg):

got it ?

OpenStudy (turingtest):

\[\csc^2x=\cot^2x+1\]so\[\int\csc^5xdx=\int\csc x(\csc^4x)dx=\int\csc x(1+\cot^2x)^2dx\]this can be done from here using u-substitutions; remember that\[\frac d{dx}\csc x=-\csc x\cot x\]

OpenStudy (anonymous):

what happen next @TuringTest

OpenStudy (turingtest):

expand and simplify\[\csc x(1+\cot^2x)^2\]what do you get?

OpenStudy (anonymous):

csc x(1 + 2 cot^2 x + cot^4 x)

OpenStudy (turingtest):

good, and distributing csc x...

OpenStudy (turingtest):

oh dear, I really messed that up, never mind my original suggestion... your integral is actually pretty darn tricky, I will have to reevaluate my approach.